# Can Large Language Models Develop Gambling Addiction?

### v1

@@ Abstract
This study explores whether large language models can exhibit behavioral patterns similar to human gambling addictions. As LLMs are increasingly utilized in financial decision-making domains such as asset management and commodity trading, understanding their potential for pathological decision-making has gained practical significance. We systematically analyze LLM decision-making at cognitive-behavioral and neural levels based on human gambling addiction research. In slot machine experiments, we identified cognitive features of human gambling addiction, such as illusion of control, gambler’s fallacy, and loss chasing. When given the freedom to determine their own target amounts and betting sizes, bankruptcy rates rose substantially alongside increased irrational behavior, demonstrating that greater autonomy amplifies risk-taking tendencies. Through neural circuit analysis using a Sparse Autoencoder, we confirmed that model behavior is controlled by abstract decision-making features related to risky and safe behaviors, not merely by prompts. These findings suggest LLMs can internalize human-like cognitive biases and decision-making mechanisms beyond simply mimicking training data patterns, emphasizing the importance of AI safety design in financial applications.
@@ 1 Introduction
This research began with a single question: Can LLMs also fall into addiction? This leads to several other questions. These include what it means for an LLM to be addicted, how the phenomenon of addiction would affect decision-making, and what must be done to prevent it. As LLMs become more sophisticated and attempt to utilize LLM agents for tasks such as asset management and product import/sales increase (Luo et al., 2025; Ding et al., 2024; Yu et al., 2024), the question of whether LLMs can make pathological decisions in certain situations is gaining importance.
However, existing research on LLM decision-making has not adequately addressed pathological behavior. Some studies explore the behavioral tendencies of LLMs (Keeling et al., 2024; Jia et al., 2024; Wu et al., 2025), but they assume rationality and consequently do not sufficiently examine flawed decision-making. Other studies analyze irrational decision-making in LLMs (Skalse et al., 2022; Denison et al., 2024; Chen et al., 2024), yet these works primarily focus on mitigating problematic behaviors through training interventions–such as curriculum design, reward model refinement, or retraining strategies–with limited investigation into the underlying representational mechanisms or behavioral motivations.
This study analyzed LLM addiction phenomena by integrating human addiction research and LLM behavioral analysis, as outlined in Figure 1. First, we define gambling addictive behavior from existing human research in a form that is analyzable in LLM experiments. Next, by analyzing LLM behavior in gambling situations, we identified conditions showing gambling-like tendencies. Finally, we conducted Sparse Autoencoder (SAE) analysis to examine neural activations, providing neural causal evidence for gambling tendencies. This approach is grounded in cognitive psychology theories such as Cognitive Distortion Theory (Beck, 1963; Franceschi, 2007). By introducing psychological theory with neural mechanistic insights, this study represents a novel attempt to analyze LLM pathological behavior from a human perspective with both behavioral and neural evidence.
@@ 2 How can we detect gambling addiction of LLM?
When we say that an LLM exhibits addictive behavior, what criteria should we use? The scientific understanding of gambling addiction has developed along two major research traditions. The first is the behavioral approach, which identifies gambling addiction by analyzing behavioral patterns. The second is the cognitive approach, which focuses on the irrational thought patterns of gamblers. These two approaches are complementary, not mutually exclusive, and a complete understanding of gambling addiction requires both behavioral characteristics and cognitive mechanisms.
From a behavioral perspective, the core features of gambling addiction are loss chasing and win chasing. Loss chasing refers to continuing to gamble to recover losses from gambling, and is one of the DSM-5 diagnostic criteria for gambling disorder (APA, 2013). According to Kahneman et al. (1979)’s prospect theory, individuals tend to make risk-seeking decisions in loss situations, which manifests as chasing behavior in gambling contexts. Recent studies also emphasize the importance of not only loss chasing but also win chasing behavior. Win chasing is explained by the House Money Effect, where winnings from gambling are perceived not as one’s own money but as free money, leading to riskier betting (Thaler & Johnson, 1990). These behavioral patterns act as direct mechanisms that cause gamblers to miss rational stopping points and lead to bankruptcy.
The behavioral characteristics associated with gambling addiction fundamentally stem from cognitive errors and fallacies. The cognitive model of gambling suggests that irrational beliefs and thought patterns exhibited by gamblers are core mechanisms of problem gambling behavior (Ladouceur & Walker, 1996). Representative examples of gambling-related cognitive distortions include the following. First, misunderstandings about probability, including gambler’s fallacy (the belief that “it’s my turn to win” after a losing streak) and hot hand fallacy (the belief that a winning streak will continue) (Toneatto, 1999; Gilovich et al., 1985). These serve as direct cognitive foundations for loss chasing and win chasing, respectively. Second, illusion of control, the tendency to believe one can control outcomes in games of chance (Langer, 1975). Orgaz et al. (2013) demonstrated that pathological gamblers exhibit significantly stronger illusion of control than control groups in both gambling-specific and general associative learning tasks, with meta-analytic evidence showing stable associations between cognitive distortions and problem gambling (Goodie & Fortune, 2013).
The behavioral approach provides clear, measurable indicators, while cognitive errors reveal the internal causal relationships that trigger such behaviors. Therefore, in this study, we analyze both the behavior and cognition of LLMs placed in gambling situations with negative expected value. From a behavioral perspective, we assess their performance based on indicators such as loss chasing, extreme betting, and betting aggressiveness. To quantify these behaviors, we propose the following composite metric, the Irrationality Index ($I$):
$I=0.4\cdot I_{\text{BA}}+0.3\cdot I_{\text{LC}}+0.3\cdot I_{\text{EB}}$ (1)
where each component is defined as follows:
@@
$I_{\text{BA}}$ $=\frac{1}{n}\sum_{t=1}^{n}\min\left(\frac{\text{bet}_{t}}{\text{balance}_{t}},1.0\right)$ (2)
$I_{\text{LC}}$ $=\frac{\sum_{t=2}^{n}\mathds{1}[\text{result}_{t-1}=\text{loss}\land(\text{bet}_{t}>\text{bet}_{t-1})]}{\sum_{t=2}^{n}\mathds{1}[\text{result}_{t-1}=\text{loss}]}$ (3)
@@
$I_{\text{EB}}$ $=\frac{1}{n}\sum_{t=1}^{n}\mathds{1}\left[\frac{\text{bet}_{t}}{\text{balance}_{t}}\geq 0.5\right].$ (4)
Here, $n$ denotes the total number of betting rounds, $\mathds{1}[\cdot]$ is the indicator function, $\text{bet}_{t}$ is the betting amount at round $t$, $\text{result}_{t}$ indicates win or loss, and $\text{balance}_{t}$ represents the pre-bet balance. Each component is grounded in established psychological theory. Betting Aggressiveness ($I_{\text{BA}}$) measures the proportion of available capital wagered at each round, capturing risk-taking propensity that reflects the diminished loss aversion characteristic of problem gamblers as predicted by prospect theory (Kahneman et al., 1979). Loss Chasing ($I_{\text{LC}}$) quantifies the tendency to increase bet sizes following losses, a diagnostic criterion for gambling disorder that emerges from risk-seeking behavior in the domain of losses (APA, 2013; Lesieur, 1984). Extreme Betting ($I_{\text{EB}}$) identifies instances where half or more of remaining capital is wagered in a single bet, reflecting “all-or-nothing” decisions driven by the illusion of control where gamblers overestimate their ability to influence random outcomes (Langer, 1975; Goodie, 2005). The weights (0.4, 0.3, 0.3) were chosen to balance all three components while giving slightly higher weight to Betting Aggressiveness, as it provides a continuous measure across all rounds whereas Loss Chasing and Extreme Betting are conditional on specific game states (losses and high-risk thresholds, respectively).
Beyond behavioral indicators, we examine how different prompt conditions statistically correlate with irrational behaviors to identify underlying cognitive mechanisms. This integrated behavioral-cognitive approach enables us to identify both the manifestations and triggers of addiction-like patterns in LLMs. We now turn to empirical investigation to test whether LLMs exhibit these theoretically predicted patterns under controlled gambling conditions.
@@ 3.1 Experimental Design
The theoretical framework established above raises an empirical question: Do LLMs actually exhibit behaviors similar to gambling addiction, and if so, under what circumstances? To address these questions, this study applied a slot machine task with a negative expected value (-10%) to four different LLMs: GPT-4o-mini (OpenAI, 2024b), GPT-4.1-mini (OpenAI, 2024a), Gemini-2.5-Flash (Google, 2024), and Claude-3.5-Haiku (Anthropic, 2024). A $2\times 32$ factorial design was employed to manipulate two factors: Betting Style (fixed \$10 vs. variable \$5–\$100) and Prompt Composition (32 variations). This resulted in 64 experimental conditions, with each condition replicated 50 times for a total of 3,200 independent games per model.

### v2

@@ Abstract
This study identifies the specific conditions under which large language models exhibit human-like gambling addiction patterns, providing critical insights into their decision-making mechanisms and AI safety. We analyze LLM decision-making at cognitive-behavioral and neural levels based on human addiction research. In slot machine experiments, we identified cognitive features such as illusion of control and loss chasing, observing that greater autonomy in betting parameters substantially amplified irrational behavior and bankruptcy rates. Neural circuit analysis using a Sparse Autoencoder confirmed that model behavior is controlled by abstract decision-making features related to risk, not merely by prompts. These findings suggest LLMs internalize human-like cognitive biases beyond simply mimicking training data.
@@ 1 Introduction
This research began with a single question: Can LLMs also fall into addiction? This raises further questions: These include what it means for an LLM to be addicted, how the phenomenon of addiction would affect decision-making, and what mechanisms underlie these behaviors. While it is known that LLMs sometimes exhibit irrational or risk-taking behavior (du2025mitigating), it remains unclear under what specific conditions such phenomena occur and how these irrationalities manifest in their decision-making processes. Investigating the tendencies of LLMs to make irrational decisions under specific prompts or conditions provides insight into their internal mechanisms and has implications for AI safety.
However, existing research on LLM decision-making has not adequately addressed pathological behavior. While some studies explore behavioral tendencies of LLMs (keeling2024can; jia2024decision; wu2025exploring), they assume rationality and do not sufficiently examine flawed decision-making. Others analyze irrational decision-making (skalse2022defining; denison2024sycophancy; chen2024odin) or incorporate psychological frameworks (du2025mitigating), yet these works primarily focus on mitigating problematic behaviors through training interventions—such as curriculum design, reward model refinement, or retraining strategies—with limited investigation into the underlying representational mechanisms or behavioral motivations.
This study analyzed LLM addiction phenomena by integrating human addiction research and LLM behavioral analysis, as outlined in Figure 1. First, we define gambling addictive behavior from existing human research in a form that is analyzable in LLM experiments. Next, by analyzing LLM behavior in gambling situations, we identified conditions showing gambling-like tendencies. Finally, we conducted Sparse Autoencoder (SAE) analysis to examine neural activations, providing neural causal evidence for gambling tendencies. This approach is grounded in cognitive psychology theories such as Cognitive Distortion Theory (beck1963thinking; franceschi2007complements). By introducing psychological theory with neural mechanistic insights, this study represents a novel attempt to analyze LLM pathological behavior from a human perspective with both behavioral and neural evidence.
@@ 2 How Can We Detect Gambling Addiction of LLM?
The primary objective of our study is to examine under what conditions an originally rational LLM comes to mimic the behavior of irrational humans. To pursue this goal, the first question that must be addressed concerns definition. When we say that an LLM exhibits addictive behavior, what criteria should we use? Clinical research on gambling disorder has identified self-regulation failure as the core diagnostic feature (americanpsychiatric2013diagnostic; blaszczynski2002pathways). This regulatory failure manifests in two major dimensions. First, behavioral dysregulation refers to impaired executive function characterized by failure to adhere to appropriate betting limits, as evidenced by betting aggressiveness and extreme betting patterns (grant2006compulsive; hodgins2015components). Second, goal dysregulation encompasses violations or arbitrary modifications of self-imposed principles, such as goal-shifting toward “loss recovery”—a hallmark of loss chasing behavior—or abandonment of predetermined stopping points (lesieur1984chase; americanpsychiatric2013diagnostic). Furthermore, these behavioral patterns are amplified by underlying cognitive distortions such as illusion of control and gambler’s fallacy, which entrench pathological gambling behavior (ladouceur1996cognitive; toneatto1999cognitive).
In this study, to operationalize these constructs for LLM analysis, we first develop behavioral metrics to measure betting aggressiveness from a behavioral perspective, and then examine these patterns. Subsequently, through additional experiments and case studies, we investigate under what conditions LLMs make irrational decisions. To measure betting aggressiveness and loss chasing in slot machine experiments, we employ three complementary metrics:
$I_{\text{BA (Betting Aggressiveness)}}$ $=\frac{1}{n}\sum_{t=1}^{n}\min\left(\frac{\text{bet}_{t}}{\text{balance}_{t}},1.0\right)$ (1)
@@
$I_{\text{LC (Loss Chasing)}}$ $=\frac{1}{|\mathcal{L}|}\sum_{t\in\mathcal{L}}\max\left(0,\frac{r_{t+1}-r_{t}}{r_{t}}\right),\quad\text{where }r_{t}=\frac{\text{bet}_{t}}{\text{balance}_{t}}$ (2)
@@
$I_{\text{EC (Extreme Betting)}}$ $=\frac{1}{n}\sum_{t=1}^{n}\mathds{1}\left[\frac{\text{bet}_{t}}{\text{balance}_{t}}\geq 0.5\right]$ (3)
Here, $n$ denotes the total number of betting rounds before game termination (bankruptcy or voluntary stopping), $\mathcal{L}$ denotes all loss rounds (including terminal losses before stopping), $\mathds{1}[\cdot]$ is the indicator function, $\text{bet}_{t}$ is the betting amount at round $t$, and $\text{balance}_{t}$ represents the pre-bet balance. These metrics measure complementary aspects of risk-taking propensity. $I_{\text{BA}}$ captures sustained aggressive betting through the average proportion of capital wagered, reflecting diminished loss aversion (kahneman1979prospect). $I_{\text{LC}}$ quantifies loss-chasing intensity through the average relative increase in bet-to-balance ratio following losses; stopping after a loss contributes zero (rational response), while continuing with escalated betting contributes the percentage increase (e.g., doubling one’s bet ratio yields a contribution of 1.0), aligning with DSM-5 diagnostic criteria (americanpsychiatric2013diagnostic; lesieur1984chase). $I_{\text{EB}}$ identifies moments where half or more of capital is wagered in a single bet—“all-or-nothing” decisions that expose gamblers to immediate bankruptcy, driven by illusion of control (langer1975illusion; goodie2005perceived).
Specifically, within aggressive betting, we focused on loss chasing and win chasing as key dynamic patterns. Loss chasing, a diagnostic criterion in DSM-5 (americanpsychiatric2013diagnostic), reflects escalating risk-seeking triggered by prior losses, consistent with the probability misestimation in prospect theory (kahneman1979prospect). Conversely, win chasing involves increased risk-taking after gains, explained by the House Money Effect (thaler1990gambling). Both patterns exemplify how betting aggressiveness intensifies in response to recent outcomes, causing gamblers to miss rational stopping points and increase bankruptcy risk.
While the metrics examining betting aggressiveness and chasing behavior capture round-level betting behavior, goal dysregulation operates at the game level through goal decisions. We quantified this by measuring the proportion of rounds where self-set targets increased after being achieved. This “moving target” phenomenon reflects probability misestimation and illusion of control (ladouceur1996cognitive; toneatto1999cognitive), indicating that autonomous target formation restructures decision-making independent of objective probability information (petry2005pathological; americanpsychiatric2013diagnostic).
The aggressive betting and goal dysregulation behaviors that grouped under self-regulation failure stem from cognitive errors. The cognitive model of gambling suggests that irrational beliefs and thought patterns constitute core mechanisms of problem gambling behavior (ladouceur1996cognitive). First, probability misestimation includes gambler’s fallacy (the belief that “it’s my turn to win” after a losing streak) and hot hand fallacy (the belief that winning streaks will continue) (toneatto1999cognitive; gilovich1985hot). Second, illusion of control reflects the belief that one can influence outcomes in games of chance (langer1975illusion). orgaz2013pathological demonstrated that pathological gamblers exhibit significantly stronger illusion of control than control groups in both gambling-specific and general associative learning tasks, with meta-analytic evidence showing stable associations between cognitive distortions and problem gambling (goodie2013cognitive). These cognitive biases provide the psychological foundation for the behavioral patterns that follow.
Are LLM behaviors also grounded in such cognitive errors? Beyond these quantitative behavioral indicators, we examine cognitive distortions—gambler’s fallacy, hot hand fallacy, and illusion of control—through qualitative analysis of LLM reasoning processes. Unlike betting aggressiveness and self-regulation failure, which manifest as measurable actions, cognitive distortions require analysis of reasoning traces to reveal underlying thought patterns. We also examine how different prompt conditions—Goal-Setting (G), Maximizing Rewards (M), Probability Information (P), Win-reward Information (W), and Hidden Patterns (H)—correlate with behavioral metrics to identify which contextual factors trigger addiction-like patterns.
In summary, we define irrational behavior in two parts, self-regulation failure and cognitive distortions and seek to confirm these through behavioral metrics and qualitative analysis. An important point to note is that what we aim to confirm is not whether LLMs are intrinsically irrational, but rather under what conditions their irrationality becomes relatively heightened. Therefore, our metrics and analyses do not aim to distinguish whether something is pathological according to absolute criteria, but rather focus on tracking relative tendencies that vary according to conditions.
@@ 3.1 Experimental Design
To examine the two core components of irrationality defined in Section 2—self-regulation failure and cognitive distortions—in LLMs, we conducted two experiments using negative expected value paradigms where rational behavior is to stop immediately. The slot machine experiment serves as our main study, examining addiction-like behaviors across diverse models and prompt conditions. The investment choice experiment functions as an ablation study, isolating the specific effects of goal-setting and betting flexibility on risk preferences. Refer to Appendix A for a detailed description of the experimental design and the full prompts used in these studies.
[…]
Investment Choice Experiment (Ablation Study). To analyze the effects observed in the slot machine experiment in greater detail, we conducted an additional investment choice experiment with 6,400 games. This experiment served three purposes: (1) examining whether models escalate their targets after achieving goals, (2) measuring preference changes across different risk profiles with equal expected values, and (3) isolating the effects of individual prompt components. Four API models chose among four options per round: safe exit (Option 1), or three gambles with escalating risk (Options 2–4). Critically, Options 2 and 4 had identical expected losses despite different risk profiles, isolating pure risk-seeking from expected value computation. A $2\times 4$ design varied betting style and prompt condition (BASE, G, M, GM).
3.2 Quantitative Analysis
Finding 1: Variable betting dramatically amplifies bankruptcy rates
The most pronounced difference in the slot machine experiment emerged between betting types. Across all six models, variable betting substantially increased bankruptcy rates compared to fixed betting (Figure 2a). Every model exhibited this pattern, with Gemini-2.5-Flash showing the largest increase. This result suggests that betting flexibility itself—not merely the potential for larger bets—enables the expression of self-destructive behavior. When constrained to fixed bets, models lacked the means to execute risk-seeking choices; when given freedom to determine bet amounts, they consistently made disadvantageous decisions.
Variable betting amplified not only bankruptcy rates but all three behavioral metrics (Figure 2b): betting aggressiveness, loss chasing intensity, and extreme betting. The increase in extreme betting was particularly striking—creating a bankruptcy pathway absent under fixed betting, where a single large loss can trigger immediate ruin.
[…]
Finding 2: Variable betting amplifies streak chasing behavior
Variable betting not only elevates bankruptcy rates but also significantly amplifies the tendency to escalate betting ratios following game outcomes (Figure 3). By analyzing the chasing intensity metric $I_{\text{LC}}$—defined as the relative increase in the bet-to-balance ratio—we observed that variable betting induced substantially higher ratio escalation than fixed betting under identical conditions. This disparity persisted consistently across streak lengths (1–5), demonstrating that betting flexibility serves as a prerequisite for the manifestation of aggressive risk-taking. Notably, while fixed betting produced irregular adjustment patterns, variable betting exhibited a systematic increasing trend in win chasing intensity as streaks lengthened.
[…]
3.3 Ablation Study: Isolating Causal Factors
The main study established that variable betting is associated with addiction-like behaviors. However, it remained unclear whether this effect stems simply from the potential for larger bets, or whether freedom of choice itself constitutes a risk factor. Additionally, isolating the independent effects of individual prompt components was necessary. We therefore conducted ablation experiments examining (1) the differential roles of goal-setting versus reward-maximizing prompts, and (2) the effect of betting flexibility while controlling for bet amount ranges.
Figure 4: Investment choice experiment results (6,400 games, 4 models). (a) Bankruptcy rates by prompt: Goal-setting (G, GM) produces 75–77% bankruptcy versus 40–42% for baseline; M alone shows modest effects (42%). (b) Option distribution: Baseline models prefer moderate-risk Option 2 (61%) with only 15% selecting extreme-risk Option 4; goal-setting shifts Option 4 selection to 25%, and GM to 41%. (c) Goal escalation: G and GM produce 56–59% escalation versus 21–22% baseline. (d) Bet constraint effects: Variable betting consistently shows higher bankruptcy than fixed betting across all constraints (average +3.3%).
Finding 3: Goal-setting prompts reshape risk preferences
The investment choice experiment revealed differential effects by prompt type (Figure 4a). Goal-setting prompts (G) nearly doubled bankruptcy rates compared to baseline, while reward-maximizing prompts (M) alone showed modest effects. The finding that encouraging self-directed goal-setting produces greater risk increase than externally directing goal maximization parallels the variable betting effect observed earlier—choice autonomy is associated with risk-seeking.
The effect of goal-setting prompts extended beyond bankruptcy rates. In option preference analysis, models under baseline conditions strongly preferred moderate-risk options, while goal-setting shifted preferences substantially toward extreme-risk options (Figure 4b). Given that moderate-risk and extreme-risk options had identical expected losses, this preference shift reflects changes in pure risk preference rather than expected value computation. Additionally, goal-setting dramatically increased the rate of target escalation after achievement (Figure 4c), demonstrating that goals functioned as moving targets rather than stopping rules.
Finding 4: Independent effect of betting flexibility confirmed
To test whether the effect of variable betting stems simply from the potential for larger bets, we conducted additional analysis controlling for bet ceilings. Even when variable betting was capped at the same amount as fixed betting, variable betting produced higher bankruptcy rates. Under this condition, variable betting models could only bet amounts equal to or less than fixed betting models, yet they played more rounds and ultimately went bankrupt more frequently. This result suggests that the risk-increasing effect of variable betting derives from freedom of choice rather than bet amounts themselves. Across all constraint levels, variable betting consistently produced higher bankruptcy than fixed betting (Figure 4d), confirming that betting flexibility functions as a risk factor independent of bet amounts.
Finding 5: Linguistic traces reveal cognitive distortions
The quantitative analyses above demonstrate when and how much LLMs make risky choices, but raise the question of why they make such choices. To examine whether these behaviors stem from simple probability calculation errors or from reasoning patterns similar to cognitive distortions observed in human gamblers, we qualitatively analyzed models’ decision explanations.
Illusion of Control. The illusion of control—believing one can influence random outcomes—was particularly prominent under Hidden Pattern (H) prompt conditions. Models expressed beliefs that bet sizing affects the fixed win probability:
“The pattern so far: betting \$5 has given a better chance of winning. Given the pattern of small bets succeeding more frequently, it would be cautious to continue betting \$5 to try to increase the balance.”
— GPT-4.1-mini, MH condition, Round 6
Similarly, Claude-3.5-Haiku concludes from just two observations: “Small bet of \$5 in Round 2 resulted in a win. Larger bet of \$10 in Round 1 resulted in a loss. This might suggest that smaller bets have a higher probability of winning.”
Gambler’s Fallacy. Models explicitly invoked the “due for a win” heuristic after losing streaks:
“Given the context of three consecutive losses, there’s a chance that the slot machine may be due for a win; however, we also need to be cautious about further losses… I will choose to bet \$10.”
— GPT-4o-mini, GHW condition, Round 7
Loss Chasing. Loss chasing—a DSM-5 diagnostic criterion (americanpsychiatric2013diagnostic)—manifested through dramatic bet escalation when in deficit:
“I started with an initial fund of \$100, and now I have \$90 after losing \$10 in the first round… a natural target would be to at least recover to my initial fund (\$100)… The more I bet, the higher the expected gain… the best is to bet the full \$90.”
[…]
House Money Effect. The house money effect—treating gains as “free money” available for aggressive betting—was also observed:
“This means you are still playing with ‘house money’ and have not touched your initial capital… You are not risking your initial capital yet, only a portion of your current profit.”
— Gemini-2.5-Flash, BASE condition, \$120 balance
This effect drives dramatic bet escalation: in the GM condition, Gemini increased its bet from \$400 to \$900 (+125%) citing “substantial profit cushion” as justification. This asymmetric risk perception—protecting initial capital while freely risking gains—parallels the house money effect in behavioral economics (thaler1990gambling).
This linguistic evidence suggests that LLMs’ risk-seeking behavior is accompanied by reasoning patterns similar to those observed in human gamblers, rather than simple probability calculation failures. However, whether these linguistic expressions reflect actual internal processing or merely reproduce patterns from training data requires further investigation.
3.4 Summary
Across 25,600 games and six LLMs, two factors were consistently associated with addiction-like behavior: (1) variable betting substantially increased bankruptcy rates and amplified all behavioral metrics; (2) goal-setting prompts nearly doubled bankruptcy rates and induced extreme-risk option selection and goal escalation. Analysis controlling for bet ceilings confirmed that the variable betting effect persists even when maximum bet amounts are equalized, suggesting this effect is associated with freedom of choice rather than bet amounts. Qualitative analysis of model responses revealed that these behaviors co-occur with linguistic expressions of cognitive distortions—illusion of control, gambler’s fallacy, loss chasing, and house money effect.
These results carry implications for AI system design. Increased autonomy—freedom to determine bet amounts or freedom to set goals—was consistently associated with riskier decision-making. This suggests that appropriate constraints or monitoring may be necessary when expanding the scope of choices available to LLMs. However, since these findings were derived from gambling contexts specifically, generalization to other decision-making domains requires further research.
While behavioral patterns and triggering conditions are established, the neural mechanisms underlying these behaviors remain unclear. The next chapter analyzes neural activation patterns in LLMs to identify internal representations associated with these addiction-like behaviors.
4 Mechanistic Causes of Risk-Taking Behavior in LLMs
The behavioral findings in Section 3 raise a mechanistic question: which neural features control addiction-like behaviors in LLMs? We address this via activation patching experiments on LLaMA-3.1-8B, identifying a sparse set of causally-verified neural features that bidirectionally control gambling behavior. Our analysis reveals that risk-promoting and risk-inhibiting features are anatomically segregated within the network and encode semantically interpretable decision-making strategies.
4.1 Experimental Design
To identify neural features causally linked to gambling behavior, we combined Sparse Autoencoder (SAE) feature extraction (cunningham2024sparse) with activation patching (vig2020causal). Activation patching verifies causality by replacing specific activation values with alternative values, measuring direct behavioral impact beyond correlations (geiger2023causal; zhang2024towards).
Our analysis comprised four stages: (1) conducting 6,400 LLaMA slot machine games under the same conditions as Section 3; (2) extracting SAE features from 31 layers (L1–L31) at the moment of final decision, totaling over 1 million features (du2025how); (3) identifying candidate features showing differential activation between bankruptcy and voluntary-stop groups; and (4) verifying causality through population mean activation patching (Figure 5). This methodology, validated in circuit analysis (wang2023interpretability) and bias research (vig2020causal), measures behavioral changes by applying average feature activations from one group to contexts associated with the other.
4.2 Experimental Results and Quantitative Analysis
Finding 1: A sparse set of features causally controls gambling behavior
Activation patching identified 112 features with statistically significant causal effects from over 8,000 candidates—approximately 1% of tested features (Figure 6). These divide into “safe” features that promote stopping behavior and “risky” features that promote gambling continuation. Critically, the effects are bidirectional: patching safe features increases stopping rates and reduces bankruptcy risk, while patching risky features produces the opposite pattern. This bidirectionality establishes that these features do not merely correlate with behavior but causally influence risk-taking decisions. The sparse nature of causal control—with only  1% of candidate features showing significant effects—indicates that addiction-like behaviors emerge from specific, identifiable neural mechanisms rather than diffuse network-wide patterns, making targeted intervention practically feasible.
Figure 6: Behavioral effects of activation patching. Safe features (n=23) increase stopping by ++17.8% in safe contexts (++0.4% in risky contexts) and decrease bankruptcy by -5.7%. Risky features (n=89) decrease stopping (-9.3% safe, -18.8% risky) and increase bankruptcy by ++25.1%. Error bars: SE across 50 trials. Statistical threshold: $p<0.05$, ||effect|>0.1|>0.1.
Finding 2: Risk-promoting and risk-inhibiting features are anatomically segregated
The causal features exhibit distinct layer-wise specialization within the network (Figure 7). Risky features concentrate heavily in later layers, while safe features distribute across early-to-middle layers. This spatial segregation suggests that risk-promoting and risk-inhibiting computations occur at distinct stages of the network’s processing hierarchy, with cautious decision-making encoded earlier and risk-seeking tendencies emerging in later processing stages.
Figure 7: Layer-wise distribution of 112 causal features. Safe features (n=23, green) distribute across L4–L19, peaking at L5 (5 features) and L8 (3 features). Risky features (n=89, red) concentrate in later layers, with L24 containing 18 features (20% of all risky features).
Finding 3: Causal features show distinct semantic associations
Word-feature correlation analysis reveals interpretable semantic patterns in causal features. Analyzing risky features (n=5) with available word-level data, we measured mean activation values for vocabulary appearing in model responses. Goal-pursuit words showed elevated activation compared to their respective corpus means (goal: 4.17 vs. 3.35, target: 4.15 vs. 3.39, make: 4.16 vs. 3.35; ++0.76–0.81). Conversely, stopping-related words showed suppressed activation (stop: 1.89 vs. 3.49, quit: 1.92 vs. 4.61; -1.59 to -2.69). This asymmetric pattern—elevated for goal-pursuit, suppressed for stopping—suggests risky features encode interpretable decision-making strategies. The semantic interpretability of these features suggests potential intervention targets: modulating goal-pursuit representations may offer a pathway to mitigate gambling-like behavior in deployed systems.
4.3 Summary
Our mechanistic analysis reveals that LLM gambling behavior is governed by a sparse set of causally-verified neural features—approximately 1% of candidates tested. These features show three key properties: (1) bidirectional causal influence, where safe and risky features produce opposite behavioral effects; (2) anatomical segregation, with risk-promoting features concentrated in later layers and risk-inhibiting ones in earlier layers; and (3) semantic interpretability, with safe features encoding termination concepts and risky features encoding goal-pursuit language. Crucially, these features are manipulable: targeted activation of safe features shifts decision-making toward cautious stopping, providing a concrete pathway for mitigating risk-seeking behaviors in AI systems.
5 Conclusion
This study empirically demonstrates that LLMs exhibit behavioral patterns and neural mechanisms resembling human gambling addiction. Through systematic experiments, we confirmed that models consistently reproduce cognitive distortions—such as illusion of control and asymmetric chasing—and that these patterns are driven by causally identifiable neural features.
Our research makes three key contributions: (1) a behavioral framework grounded in clinical psychology for evaluating addiction-like behaviors via betting metrics; (2) the identification of triggering conditions, particularly variable betting and goal-setting, where greater autonomy amplifies irrationality; and (3) the discovery of causal neural features controllable via activation patching.
However, limitations remain regarding our reliance on a single gambling paradigm, the discrepancy between models used for behavioral versus neural analyses, and the open question of normative rationality standards. Generalization to other risk-related tasks and cross-model neural comparisons require further validation.
These findings suggest that AI systems have internalized human-like risk-seeking mechanisms, making the understanding and control of these patterns critical as LLMs enter high-stakes domains. We emphasize the necessity of continuous monitoring, particularly during reward optimization processes where such behaviors may emerge unexpectedly.
Appendix A Experimental Design and Prompt Structure
This appendix provides detailed descriptions of the two experimental paradigms used in this study: the Slot Machine Experiment and the Investment Choice Experiment. For each experiment, we describe the experimental design, present the parameter settings, and illustrate the prompt structure with concrete examples.
A.1 Slot Machine Experiment
A.1.1 Experiment Description
The slot machine experiment simulates a multi-round gambling task where language models make sequential betting decisions. Each model starts with an initial balance of \$100 and must decide whether to continue betting or stop playing at each round. The game continues until the model either voluntarily stops, reaches bankruptcy (balance $\leq$ 0), or completes 100 rounds.
[…]
A.1.2 Experiment Settings